# PaintScene4D: Consistent 4D Scene Generation from Text Prompts

### v1

@@ Abstract
Recent advances in diffusion models have revolutionized 2D and 3D content creation, yet generating photorealistic dynamic 4D scenes remains a significant challenge. Existing dynamic 4D generation methods typically rely on distilling knowledge from pre-trained 3D generative models, often fine-tuned on synthetic object datasets. Consequently, the resulting scenes tend to be object-centric and lack photorealism. While text-to-video models can generate more realistic scenes with motion, they often struggle with spatial understanding and provide limited control over camera viewpoints during rendering. To address these limitations, we present PaintScene4D, a novel text-to-4D scene generation framework that departs from conventional multi-view generative models in favor of a streamlined architecture that harnesses video generative models trained on diverse real-world datasets. Our method first generates a reference video using a video generation model, and then employs a strategic camera array selection for rendering. We apply a progressive warping and inpainting technique to ensure both spatial and temporal consistency across multiple viewpoints. Finally, we optimize multi-view images using a dynamic renderer, enabling flexible camera control based on user preferences. Adopting a training-free architecture, our PaintScene4D efficiently produces realistic 4D scenes that can be viewed from arbitrary trajectories. The code will be made publicly available.
Figure 1: 4D Text-to-Scene Generation: Unlike prior methods that restrict text-to-4D generation to object-level reconstruction or text-to-video models lacking explicit camera control, our approach reconstructs full 4D scenes from a single text prompt. This leads to realistic 4D scenes that can be viewed from different trajectories, achieving via an efficient training-free architecture.
@@ 1 Introduction
Generating dynamic 3D scenes from text descriptions, known as text-to-4D scene generation, represents one of the most challenging frontiers in computer vision and graphics. While recent advances have revolutionized our ability to create static 3D content [28, 42, 29, 33, 23, 31, 34] and 2D image and videos [49, 38, 43, 39] from text, the synthesis of temporally coherent and animated 3D scenes remains a fundamental challenge. This task requires not only generating spatially consistent 3D geometry and appearance, but also producing realistic motion that adheres to real-world physics and semantic constraints – all while maintaining temporal consistency across multiple viewpoints.
@@
The complexity of 4D scene generation stems from several interconnected challenges. First, unlike static 3D generation, which only needs to ensure spatial consistency, 4D scenes must maintain both spatial and temporal coherence simultaneously. This means that any generated motion must be physically plausible and semantically meaningful, while preserving the scene’s geometric structure across time. Second, the lack of large-scale, diverse 4D scene datasets has limited the development of robust generation methods, with most existing approaches relying on object-centric data that fail to capture the rich dynamics of full scenes. Third, the computational complexity of optimizing both spatial and temporal dimensions makes it difficult to achieve high-quality results within reasonable time constraints.
@@
Current approaches to these challenges broadly fall into two categories, each with significant drawbacks. The first category extends static 3D generation methods (e.g., MVDream [42] and Zero123 [28, 41], trained on object-centric datasets like Objaverse [9]) to incorporate temporal dynamics [25, 44, 62, 56, 36, 61, 19, 2]. These methods, while effective at maintaining geometric consistency, typically struggle with generating complex motion, often producing only subtle deformations or simple translations. This limitation stems from their reliance on Score Distillation Sampling (SDS) [33] optimization, which, while effective for static content, becomes computationally intractable when scaled to temporal sequences. The second category are the Text-to-Video models (e.g., Animatediff [13], CogVideo [54]) that generate dynamic content. However, these approaches lack explicit 3D understanding, resulting in temporal inconsistencies and geometric artifacts. Neither approach adequately addresses the fundamental challenge of generating spatially and temporally coherent 4D scenes.
@@
To address these limitations, we present PaintScene4D, a novel framework that harness the strengths of both Text-to-Video generation and 3D-aware neural rendering. Our key insight is that by using video generation as an initial prior and reconstructing the 3D scene through a progressive warping and inpainting technique, we can maintain both spatial and temporal consistency while enabling complex motion generation. Specifically, our method first generates a base video using a pretrained Text-to-Video model, which provides rich motion priors. We then construct a “web of cameras” around the scene by warping video frames to nearby viewpoints and using inpainting to fill any resulting gaps. This approach allows us to build a comprehensive multi-view representation of the dynamic scene without requiring explicit 3D supervision or costly optimization procedures.
@@
The effectiveness of PaintScene4D is demonstrated through several empirical contributions. As shown in Figure 1, our method achieves state-of-the-art results in text-to-4D scene generation, producing visually compelling results that maintain both spatial and temporal consistency. The generated scenes exhibit complex motion while preserving geometric structure across multiple viewpoints. Notably, our framework reduces the computational requirements significantly, generating high-quality 4D content in approximately 3 hours on a single A100 GPU – a substantial improvement over existing methods [2, 62] that often require 10+ hours. Through extensive experiments and ablation studies, we demonstrate the superiority of our approach across various metrics, including temporal consistency, motion complexity, and rendering quality. Our method also offers unprecedented flexibility, allowing users to edit existing videos or specify custom trajectories during inference.
Our main contributions can be summarized as follows:
@@
  • A novel end-to-end framework for text-to-4D scene generation that effectively distills video generation prior to 4D-aware neural rendering.
  • A progressive warping and inpainting technique that enables the construction of spatially and temporally consistent multi-view representations.
@@
  • Comprehensive evaluation and analysis of PaintScene4D demonstrates state-of-the-art results in 4D scene generation, with significantly reduced computational requirements and enhanced camera control options.
@@ 2 Related Work
Text-to-3D Generation.   Text-to-3D generation has evolved significantly over the past decades. Initial approaches rely on rule-based systems that parse text inputs into semantic representations for scene generation using object databases [1, 8, 6]. The field has advanced substantially with the introduction of data-driven approaches that leverage multimodal datasets [7] and pretrained models like CLIP [35], enabling more sophisticated manipulation of 3D meshes [18, 11] or radiance fields [48]. This progress has led to the development of methods utilizing CLIP-based supervision for comprehensive 3D scene synthesis [17, 40], which subsequently evolves into techniques that optimize meshes and radiance fields through Score Distillation Sampling (SDS) [33, 50, 22]. The introduction of multi-view-aware diffusion models has further enhanced the quality of generated 3D structures [24, 28, 42]. Parallel developments in diffusion and transformer architectures have enabled advanced image-to-3D conversion for novel view synthesis [5, 45, 12, 29, 57, 47, 34]. These approaches primarily address object-level reconstruction.
@@
Recent advances in text-to-3D scene generation have introduced innovative approaches to address scene-level complexity. Text2Room [14] proposes a warping and inpainting methodology for mesh population and scene creation, while Text2NeRF [59] shifts away from mesh-based reconstruction to utilize radiance fields as scene generation priors. Although these methods are initially constrained to camera-centric scenes, subsequent work [60] expands capabilities to support general 3D scene generation with arbitrary 6 degree-of-freedom (DOF) camera trajectories. However, these approaches remain limited to static scenes, lacking the ability to incorporate motion, which is a crucial element for representing dynamic, real-world environments.
@@
Object-centric Text-to-4D Generation.   The extension from 3D to 4D scene generation introduces significant additional complexity. MAV3D [44] pioneers this direction by introducing a dynamic NeRF representation using HexPlane [4] and video-based SDS loss, structuring the generation process into static, dynamic, and superresolution phases. Building on this foundation, Dream-in-4D [63] employs a deformation field-based dynamic NeRF, organizing text-to-4D generation into distinct static and dynamic phases. Similarly, 4D-fy [2] introduces a hybrid representation that combines static and dynamic voxels, leveraging multiple SDS loss functions [33, 42, 30] to enable dynamic content generation. Additionally, Ling et al. [26] achieve dynamic rendering through the application of a dynamic network to 3D Gaussian Splatting (3D-GS), optimized under both video SDS and 3D Gaussian constraints.
Recent developments have focused on decomposing and controlling motion generation. TC4D [3] separates motion into global trajectories, represented by spline curves, and local object motion, with global paths being user-defined and local motion generated in segments using video-based SDS loss. Comp4D [53] employs large language models LLMs to decompose prompts into entities, generating and combining 4D objects based on LLM-derived trajectory information. These approaches primarily focus on object-level reconstruction, limiting their broader applicability. A notable departure from this trend is 4Real [58], which circumvents multi-view generative models by leveraging video generative models trained on large-scale datasets. Our work builds upon these advances while addressing their limitations, proposing a method for generating 4D scenes that captures both geometry and texture in real-world environments while offering enhanced control over camera movement and rendering capabilities.
Figure 2: Method Overview. Our approach consists of three stages. First, we initialize the 4D scene using a diffusion prior to establish scene content and motion, estimate depth maps for each video frame, and initialize camera trajectory (extrinsics) and intrinsics for subsequent warping. In the second stage, we perform sequential warping and inpainting from the first timestamp. To ensure spatial and temporal coherence, our consistent inpainting module mitigates artifacts and aligns depth maps, preventing error accumulation. Finally, the generated view-time matrix is used to renders novel views along user-defined camera trajectories, allowing for explicit camera control. $V_{0},V_{1}$ represent different view points.
3 Preliminary: 3D Gaussian Splatting
3D Gaussian Splatting (3D-GS) [21] utilizes point clouds to explicitly represent scenes, with each point modeled as a Gaussian distribution in three-dimensional space. Each Gaussian is defined by a central point $\mathcal{X}$, representing its mean, and a covariance matrix $\Sigma$, capturing its spread. The Gaussian function $G(\mathcal{X})$ for a point is represented by:
$G(\mathcal{X})=\exp\left(-\frac{1}{2}\mathcal{X}^{T}\Sigma^{-1}\mathcal{X}\right).$ (1)
To enable differentiable optimization, the covariance matrix $\Sigma$ is decomposed into a rotation matrix $\mathbf{R}$ and a scaling matrix $\mathbf{S}$ as $\Sigma=\mathbf{R}\mathbf{S}\mathbf{S}^{T}\mathbf{R}^{T}.$ Rendering novel views involves a differentiable splatting technique [55], where 3D Gaussians are projected onto a 2D image plane. As detailed in [64], the transformed covariance $\Sigma^{\prime}$ in camera coordinates is computed using the viewing transformation matrix $W$ and the Jacobian $J$ of the projection, yielding $\Sigma^{\prime}=JW\Sigma W^{T}J^{T}.$ In this setup, each Gaussian is parameterized by its position $\mathcal{X}\in\mathbb{R}^{3}$, its color defined by spherical harmonic (SH) coefficients $\mathcal{C}\in\mathbb{R}^{k}$ (where $k$ is the number of SH functions), opacity $\alpha\in\mathbb{R}$, rotation $r\in\mathbb{R}^{4}$, and scale $s\in\mathbb{R}^{3}$. The resulting pixel color and opacity are derived from the Gaussian blend defined in Equation 2, where the color $C$ from $N$ overlapping Gaussians is calculated as:
$C=\sum_{i=1}^{N}c_{i}\alpha_{i}\prod_{j=1}^{i-1}\left(1-\alpha_{j}\right),$ (2)
where $c_{i}$ and $\alpha_{i}$ represent the color and opacity of the $i$-th Gaussian point, with these values optimized per-point from the Gaussian’s SH coefficients and adjustable opacity.
@@ 4 Method
Overview. In this work, we present PaintScene4D, a novel framework designed to generate 4D dynamic scenes from textual inputs. Our approach begins with a video diffusion model that produces an initial video serving as both scene and motion reference. Using this video as input, we employ a depth estimation model to derive depth maps from each frame, allowing us to progressively construct a spatial representation of the scene. To create a comprehensive multi-camera view of the scene, we progressively warp the initial frames to new camera positions, beginning with the first frame. In this process, any regions missing due to occlusion or perspective changes are filled in using a spatially consistent inpainting method. For each subsequent frame, our approach reuses inpainted data from prior timestamps for continuity, and only filling in new, unobserved areas. Once we have constructed a network of cameras, where each camera captures all frames over time, we employ a 4D rendering algorithm to reconstruct the scene and generate novel viewpoints. This entire methodology is outlined in Figure 2.
@@ 4.1 Scene Initialization
Camera Trajectory.   To support the intended trajectory of the final rendered output, we establish a network of virtual cameras to match the user’s desired camera path. These cameras represent a structured arrangement of views that form the backbone for the construction of the 4D scene. Given that our framework incorporates warping operations, it is imperative to obtain accurate intrinsics of the camera parameters. However, the utilization of video generated by a diffusion model presents a challenge, as direct access to the intrinsic camera parameters is not available. To address this shortcoming, we employ a pre-trained model, Perspective Field [20], to calculate the intrinsic matrix based on the video frames provided.
4.2 Scene Construction w/ Progressive Inpainting
Given the absence of multi-view supervision, directly employing a single-view video $V_{0}$ and its depth maps $D_{0}$ to train a 4D radiance field can lead to issues of overfitting and geometric ambiguity. To address this, we apply a depth image-based rendering (DIBR) technique [10] to establish a network of virtual cameras around the initial view. Specifically, for each pixel $p$ in $I^{t}_{i}$ and its corresponding depth $z$ in $D^{t}_{i}$, we compute its transformed coordinates $p_{i\rightarrow j}$ and depth $z_{i\rightarrow j}$ for a neighboring viewpoint $j$ as follows:
@@ 4.2 Scene Construction w/ Progressive Inpainting
where $\mathbf{K}$, $\mathbf{P}_{i}$ and $\mathbf{P}_{j}$ are the intrinsic matrix, camera pose for view $i$ and view $j$, respectively and $I^{t}_{i}$ represents the image at the timestamp $t$ of viewpoint $i$. Following the transformation, we fill occluded or missing regions in the newly warped views with inpainting. Our experimental findings reveal that the diffusion-based prior used for inpainting yields higher quality results when the inpainted regions are larger. Therefore, for each view, we select the farthest available viewpoint with minimal overlap, warp the current frame to this viewpoint, and apply inpainting as necessary. Large occlusions are filled using a 2D diffusion-based prior, while smaller gaps are addressed with Telea-based inpainting [46].
Our warping process begins at the first timestamp, progressively warping and inpainting frames across all views before proceeding to subsequent timestamps. For the first timestamp, we start with a base view $I^{0}_{0}$, warp it to a neighboring viewpoint $I^{0}_{1}$, and inpaint any missing regions. To ensure spatial consistency, we integrate both the original ($I^{0}_{0}$) and newly warped frames ($I^{0}_{1}$) for further warping (e.g., $I^{0}_{2}$, $I^{0}_{3}$). This approach ensures that any inpainted content in $I^{0}_{1}$ is preserved in subsequent viewpoints ($I^{0}_{2}$, $I^{0}_{3}$, ….), maintaining coherence throughout the scene.
Depth Alignment.   To transform a 2D image $I$ into a 3D representation, we first estimate the depth for each pixel. Accurate integration of both new and existing content requires precise depth alignment, ensuring that similar elements in the scene, such as walls or furniture, appear at consistent depths across views. Directly projecting the predicted depth often results in abrupt transitions and geometric discontinuities due to inconsistent scale across viewpoints. To address this, we apply a depth alignment procedure inspired by Liu et al. [27], which refines the depth through scale and shift optimization. Specifically, we optimize scale $\gamma$ and shift $\beta$ parameters, $\gamma,\beta\in\mathbb{R}$, by minimizing the difference between predicted $\hat{d}$ and rendered depths ${d}$ in a least-squares sense:
@@
where the mask $m$ excludes unobserved pixels from the alignment process.
Additionally, depth estimation models may fail to accurately resolve depth at object boundaries, often yielding smooth transitions where abrupt changes are expected. This issue affects the overall warping quality, resulting in artifacts such as trailing patterns within occluded regions. To address this, we apply bilateral filtering to sharpen the depth boundaries, enhancing inpainting performance. Additional implementation details are provided in the supplementary material.
4.3 Scene Completion with Temporal Consistency

### v2

@@ Abstract
Recent advances in generative models have revolutionized 2D and 3D content creation, yet generating photorealistic 4D scenes remains a significant challenge. Existing methods typically rely on distilling knowledge from pre-trained 3D generative models, often fine-tuned on synthetic object datasets, resulting in scenes that are object-centric and lack photorealism. While text-to-video models can generate more realistic scenes with motion, they often struggle with spatial understanding and limited camera trajectory control. To address these limitations, we present PaintScene4D, a novel text-to-4D scene generation framework that departs from conventional 4D generative models in favor of a streamlined architecture that harnesses video generative models trained on diverse real-world datasets. Our model starts with the creation of a reference video and then employs a strategic camera trajectory control and a camera array selection module for novel view rendering. We introduce a progressive warping and inpainting strategy to ensure spatial and temporal consistency. Finally, we optimize novel-view videos using a dynamic renderer, enabling flexible camera control based on user preferences. Demonstrating the first training-free approach for 4D scene generation, PaintScene4D efficiently produces realistic and dynamic scenes viewed from arbitrary trajectories. The code will be made publicly available.
Figure 1: 4D Text-to-Scene Generation. Unlike prior methods that restrict text-to-4D generation to object-level reconstruction or text-to-video models lacking explicit camera control, our approach reconstructs full realistic 4D scenes that can be viewed from different trajectories, achieving via an efficient training-free architecture.
@@ 1 Introduction
Generating dynamic 3D scenes from text descriptions, known as text-to-4D scene generation, represents one of the most challenging frontiers in vision and graphics. While recent advances have revolutionized our ability to create videos [49, 38, 43, 39] and static 3D content [28, 42, 29, 33, 23, 31, 34], the synthesis of temporally coherent and animated scenes remains a fundamental challenge.
@@
The complexity of 4D scene generation stems from several interconnected challenges. First, unlike static 3D generation, 4D scenes must maintain spatial and temporal coherence simultaneously, meaning that any generated motion must be physically plausible and semantically meaningful while preserving geometric structure over time. Second, the lack of large-scale 4D scene datasets has limited the development of robust generation methods, resulting in most existing approaches relying on object-centric data without rich dynamics of scenes. Third, the computational complexity of spatial-temporal training makes it difficult to achieve high-quality results within reasonable time constraints.
@@
Current approaches to these challenges broadly fall into two categories, each with significant drawbacks. The first category extends static 3D generation methods [42, 28, 41] trained on object-centric datasets [9] to incorporate temporal dynamics [25, 44, 62, 56, 36, 61, 19, 2]. These methods, while effective at maintaining geometric consistency, struggle with complex motion and often produce only subtle deformations and translations. The second category is text-to-video (T2V) models [13, 55] that lack explicit 3D understanding, resulting in spatial inconsistencies and geometric artifacts. Both methods require significant training time, and neither adequately addresses the fundamental challenge of generating coherent spatial-temporal 4D scenes.
@@
To address these limitations, we present PaintScene4D, a novel training-free framework that harnesses the strengths of T2V generation and 4D-aware neural rendering (note that the post-processing 4D Gaussian renderer is learnable. But it requires less than an hour and, in principle, can be achieved in a training-free manner.) Our key insight is that by using video generation as an initial prior and reconstructing the 3D scene through a progressive warping and inpainting technique, we can maintain spatial-temporal consistency while enabling complex motion generation. Specifically, our method first generates a base video using a pre-trained T2V model to provide rich motion priors. We then construct a “web of cameras” around the scene by warping the frames to novel viewpoints with a minimum-overlapping viewpoint selection. We propose the progressive warping module (PWM) and the consistent inpainting module (CIM), allowing us to determine an optimal sequence for warping and inpainting and build a consistent multi-view representation of the dynamic scene, without requiring explicit 3D supervision or costly optimization.
@@
The effectiveness of PaintScene4D is demonstrated through extensive empirical contributions. As shown in Figure 1, our method achieves state-of-the-art results in text-to-4D scene generation, producing visually compelling results that maintain spatial-temporal consistency. The generated scenes exhibit complex motion while preserving geometric structure across multiple viewpoints. Notably, our framework reduces computational requirements significantly due to its training-free manner, generating high-quality 4D content in approximately 2.2 hours on a single A100 GPU, a substantial improvement over existing methods [2, 62] that often require 10+ hours. Through extensive ablation studies, we demonstrate the superiority of our approach across various metrics, including temporal consistency, motion complexity, and rendering quality. Our method also offers notable flexibility, allowing users to edit existing videos or specify custom trajectories during inference.
Our main contributions can be summarized as follows.
@@
  • We propose a novel modularized and training-free framework for text-to-4D scene generation that effectively distills video generation prior to 4D-aware neural rendering.
  • We introduce a progressive warping and inpainting technique, and demonstrate key technical designs to achieve high-quality results with the combination of video generation and inpainting methods.
@@
  • We perform a comprehensive evaluation and analysis of PaintScene4D, showing leading results in the generation of 4D scenes, with significantly reduced computational requirements and enhanced camera control options.
Figure 2: Method Overview. Our approach consists of three stages. First, we initialize the 4D scene using a diffusion prior to establish scene content and motion, estimate depth maps for each video frame, and initialize camera trajectory (extrinsics) and intrinsics for subsequent warping. In the second stage, we perform sequential warping and inpainting from the first timestamp. To ensure spatial and temporal coherence, our consistent inpainting module mitigates artifacts and aligns depth maps, preventing error accumulation. Finally, the generated view-time matrix is used to render novel views along user-defined camera trajectories, allowing for explicit camera control.
@@ 2 Related Work
Text-to-3D Generation.   Text-to-3D generation has evolved significantly over the past decades. Initial approaches rely on rule-based systems that parse text inputs into semantic representations for scene generation using object databases [1, 8, 6]. The field has advanced substantially with the introduction of data-driven approaches that leverage multimodal datasets [7] and pre-trained models like CLIP [35], enabling more sophisticated manipulation of 3D meshes [18, 11] or radiance fields [48]. This progress has led to the development of methods utilizing CLIP-based supervision for comprehensive 3D scene synthesis [17, 40], which subsequently evolves into techniques that optimize meshes and radiance fields through Score Distillation Sampling (SDS) [33, 50, 22]. The introduction of multi-view-aware diffusion models has further enhanced the quality of generated 3D structures [24, 28, 42]. Parallel developments in diffusion and transformer architectures have enabled advanced image-to-3D conversion for novel view synthesis [5, 45, 12, 29, 57, 47, 34]. These approaches primarily address object-level reconstruction.
@@
Recent advances in text-to-3D scene generation have introduced innovative approaches to address scene-level complexity. Text2Room [14] proposes a warping and inpainting methodology for scene creation, while Text2NeRF [59] shifts away from mesh-based reconstruction to utilize radiance fields as scene generation priors. Subsequent work [60] expands the capabilities to support general 3D scene generation with arbitrary 6 degree-of-freedom camera trajectories. However, these approaches remain limited to static scenes, lacking the ability to incorporate motion, a crucial element for dynamic environments.
@@
Object-centric Text-to-4D Generation.   The extension from 3D to 4D scene generation introduces significant additional complexity. MAV3D [44] pioneers this direction by introducing a dynamic neural radiation field (NeRF) representation using HexPlane [4] and a video-based SDS loss. Dream-in-4D [62] employs a dynamic NeRF, organizing text-to-4D generation into distinct static and dynamic phases. Similarly, 4D-fy [2] introduces a hybrid representation that combines static and dynamic voxels with SDS loss functions [33, 42, 30]. Additionally, AYG [26] achieves dynamic rendering through the application of a dynamic network to 3D Gaussian splatting. Recent developments have focused on decomposing and controlling motion generation. TC4D [3] separates motion into global trajectories with user-defined global paths and local motion generated in segments. Comp4D [54] employs the large language model (LLM) to decompose the prompts into entities, generating 4D objects with LLM-derived trajectories. These approaches focus on object-level reconstruction, limiting their broader applicability.
Scene-centric Text-to-4D Generation.   A notable departure from this trend is 4Real [58], which circumvents multi-view generative models by leveraging video generative models trained on large-scale datasets. VividDream [21] generates explorable 4D scenes with ambient dynamics with pre-trained video diffusion and inpainting modules. Another concurrent work CAT4D [52] generates 4D dynamic scenes from monocular videos by a pre-trained multi-view video diffusion model and a deformable 3D Gaussian representation for novel view synthesis. Our approach differs from previous methods by introducing a training-free framework, generating 4D scenes that accurately capture both the geometric structure of real-world environments while providing greater control over camera movement and rendering. Additionally, our method enables the rendering of a 4D scene in just 2 to 3 hours, making it computationally efficient and practical for real-world applications.
@@ 3 Method
Overview. In this work, we present PaintScene4D, a novel framework designed to generate 4D dynamic scenes from textual inputs. Our approach begins with an initial diffusion-generated video that serves as both a scene and a motion reference. Using this video as input, we employ a depth estimation model to derive depth maps from each frame, allowing us to progressively construct a spatial representation of the scene. To create a comprehensive multi-camera view of the scene, we develop a progressive warping module (PWM) where regions missing due to occlusion or perspective changes during the warping process are progressively filled in using a spatially consistent inpainting module (CIM). For each subsequent frame, our approach reuses inpainted data from prior timestamps for continuity and only fills in new, unobserved areas. Once we have constructed a network of cameras, where each camera captures all frames over time, we employ a 4D rendering algorithm to reconstruct the scene and generate novel viewpoints. This entire methodology is described in Figure 2.
Figure 3: Gallery of Results. PaintScene4D successfully generates 4D scenes that maintain view- and temporal-coherence. The horizontal axis represents the time; the vertical axis represents different viewpoints. More visualizations are provided in the supplementary materials.
@@ 3.1 Scene Initialization
Camera Trajectory.   To support the intended trajectory of the final rendered output, we establish a network of virtual cameras to match the user’s desired camera path. These cameras represent a structured arrangement of views that form the backbone for the construction of the 4D scene. Given that our framework incorporates warping operations, it is imperative to obtain accurate intrinsics of the camera parameters in the generated videos. To address this, we employ a pre-trained model, Perspective Field [20], to calculate the intrinsic matrix based on the video frames provided.
3.2 Progressive Warping Module (PWM)
Given the absence of multi-view supervision, directly employing a single-view video $V_{0}$ and its depth maps $D_{0}$ to train a 4D radiance field can lead to issues of overfitting and geometric ambiguity. To address this, we apply a depth image-based rendering technique (DIBR [10]) to establish a network of virtual cameras around the initial view. Specifically, for each pixel $p$ in $I^{t}_{i}$ and its corresponding depth $z$ in $D^{t}_{i}$, we compute its transformed coordinates $p_{i\rightarrow j}$ and depth $z_{i\rightarrow j}$ for a neighboring viewpoint $j$ as follows:
@@ 3.2 Progressive Warping Module (PWM)
where $\mathbf{K}$, $\mathbf{P}_{i}$ and $\mathbf{P}_{j}$ are the intrinsic matrix, camera pose for view $i$ and view $j$, respectively, and $I^{t}_{i}$ represents the image at timestamp $t$ of viewpoint $i$. Following the transformation, we fill the missing or occluded regions in the newly warped views with inpainting. Our experiments reveal that the inpainting diffusion-based prior yields higher-quality results when the inpainted regions are larger. Therefore, for each view, we select the farthest available viewpoint with minimal overlap, warp the current frame to this viewpoint, and apply inpainting as necessary. Large occlusions are filled using a 2D diffusion-based prior, while smaller gaps are addressed with Telea-based inpainting [46]. Refer to Sec. 5 for quantitative and qualitative ablation studies on this.
Our warping process begins at the first timestamp, progressively warping and inpainting frames across all views before proceeding to subsequent timestamps. For the first timestamp, we start with a base view $I^{0}_{0}$, warp it to a neighboring viewpoint $I^{0}_{1}$, and paint any missing regions. To ensure spatial consistency, we integrate both the original ($I^{0}_{0}$) and newly warped frames ($I^{0}_{1}$) for further warping (e.g., $I^{0}_{2}$, $I^{0}_{3}$). This approach ensures that any inpainted content in $I^{0}_{1}$ is preserved in subsequent viewpoints ($I^{0}_{2}$, $I^{0}_{3}$, etc.), maintaining coherence throughout the scene.
Depth Alignment.   To transform a 2D image $I$ into a 3D representation, we first estimate the depth for each pixel. Accurate integration of both new and existing content requires precise depth alignment, ensuring that similar elements in the scene, such as walls or furniture, appear at consistent depths across views. Directly projecting the predicted depth often results in abrupt transitions and geometric discontinuities due to inconsistent scale across viewpoints. To address this, we apply a depth alignment procedure inspired by Liu et al. [27], which refines the depth through scale and shift optimization. Specifically, we optimize scale $\gamma$ and shift $\beta$ parameters $\gamma,\beta\in\mathbb{R}$ by minimizing the difference between the predicted depth $\hat{d}$ and the rendered depth ${d}$ in the least-squares sense.
@@
Where mask $m$ excludes unobserved pixels from the alignment. Additionally, depth estimation models may fail to accurately resolve depth at object boundaries, often yielding smooth transitions where abrupt changes are expected. This issue affects the overall warping quality, resulting in artifacts such as trailing patterns within occluded regions. To address this, we apply bilateral filtering to sharpen the depth boundaries, enhancing inpainting performance. Additional details are provided in the supplementary material.
3.3 Consistent Inpainting Module (CIM)